The following information for a company is given:

Fixed cost per month

$2,500

Unit selling price

$   100

Variable cost as a percentage of sales

60%

What amount of annual sales must the company achieve to break even?

A- $100,000

B- $75,000

C-$50,000

D-$30,000
Answer #1

B $ 75,000

Break even point is that amount of sales where the company has no profit no loss.

Calculation of Contribution per unit:

Unit Selling price = $ 100

Less: Variable cost (60%) = 60

Contribution per unit = $ 40

Calculation of Break Even:

Fixed cost per month = $ 2,500

Annual Fixed cost (2500*12) = $ 30,000     

Contribution per unit = $ 40

Break Even Point (unit) = Annual Fixed cost / contribution per unit

= 30,000 / 40 = 750 unit

Break Even point ($) = 750*100 = $ 75,000

Company should achive annual sales of $ 75,000 to Break even.